OpenStudy (anonymous):

2xy/3x+y What is answer when x limit zero & y limit zero ?????

OpenStudy (anonymous):

\[\lim_{x,y \to 0}\frac{2xy}{3x+y}\] divide by xy you get \[\lim_{x,y \to 0}\frac{2}{\frac{3}{y}+\frac{1}{x}}\]

OpenStudy (anonymous):

we can now evaluate \[\lim_{y \to 0}\frac{3}{y}=\infty\] and \[\lim_{x \to 0}\frac{1}{y}=\infty\] hence we have \[\lim_{x,y \to 0}\frac{2xy}{3x+y}=\frac{2}{\infty}=0\]

OpenStudy (anonymous):

thank u .....is it indeterminate form???

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

yes, jonask ?

OpenStudy (anonymous):

underterminate shud mean that we cant decide about the limit right for example \[\infty^\infty\] but here we know the limit is 0

OpenStudy (anonymous):

so its derterminate

OpenStudy (anonymous):

thanks lot

hartnn (hartnn):

and hey, lahiru, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

\[\Huge \color{purple} { WELCOME ,LAHIRU} !!!\]

OpenStudy (zarkon):

what if you come along the path \[y=-\sin(3x)\]

OpenStudy (zarkon):

\[y=-3\sin(x)\] is a little nicer.

OpenStudy (zarkon):

Just to be clear...I'm saying the limit does not exist.